OpenStudy (anonymous):

At a school carnival, the diameter of the mat of a trampoline is 10 feet and the diameter of its metal frame is 12 feet. What is the length, in feet, of the metal frame that surrounds the trampoline? Use 3.14 for π and round your answer to the nearest tenth. 15.7 feet 18.9 feet 31.4 feet 37.7 feet

OpenStudy (anonymous):

What is this asking for

OpenStudy (campbell_st):

you are being asked to find the circumference of the metal frame, which has a diameter of 12 ft. use the formula \[C = \pi \times d\] just substitute the information you have been given \[\pi = 3.14~~and~~~d = 12\] hope it helps